OpenStudy (anonymous):

slope=1 y-intercept = 0 a. y = –x + 1 c. y = –x b. y = x d. y = x + 1 help

OpenStudy (anonymous):

D

OpenStudy (lalaly):

y=mx+b where m is the slope and b is the y intercept

OpenStudy (austinl):

y=mx+b Slope intercept m=slope b=y intercept. And @lalaly beat me to it.

OpenStudy (anonymous):

The equation of the line is : \(y=ax+b\) where \(a\) is the slope : \(1\) is a slope means \(a=1\), and \(y-intercept=0\) means when \(x=0\) the line intercept the y axis in \(0\), mathematically : \(0=1(0)+b <===> b=0\) Finally : \(y=x\) because \(a=1\) and \(b=0\)

OpenStudy (texaschic101):

I agree...the answer would be y = x + 0, but you can just leave the 0 off and make it y = x